OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundreth. 5 = ln(x + 1)

OpenStudy (anonymous):

x = \[x = e^{5} + 1\] Wanna explanation ?

OpenStudy (anonymous):

these are my options. 53.60 54.60 146.41 147.41

Nnesha (nnesha):

to cancel out ln you have to take e both side for example \[\huge\rm \ln x \] take e \[\large\rm e^{\ln} x \] e can cancel out ln so \[\large\rm \cancel{e^{\ln}} x = x\]

Nnesha (nnesha):

take e both side and let me know what you get after that

OpenStudy (anonymous):

Where did you get e from?

Nnesha (nnesha):

i took e both side just like we take square both side to cancel out square root

Nnesha (nnesha):